Solve
a) x^3+8x^2+17x+10=0

Can you please help me out? Thanks so much in advance:)
Can you please show the steps it would really help me understand:)

Answer by lwsshak3(11628)

Solve
a) f(x)=x^3+8x^2+17x+10=0
Use rational roots theorem to solve:
....0...|......1.......8........17.......10...........
....1...|......1......9.........26.......36 upper boundary (all numbers positive)
..-1...|......1......7.........10.....-10 (-1 is a root)
f(x)=(x+1)(x^2+7x+10=0
f(x)=(x+1)(x+2)(x+5)=0
roots: x=-5, -2,-1